# TinyMyo: a Tiny Foundation Model for Flexible EMG Signal Processing at the Edge

### v1

@@ 3.1.4 Pre-training Datasets
For model pre-training, we employ three publicly available datasets, with a summary provided in Table 2.
Table 2: Pre-training EMG datasets.
Dataset #Sub $f_{s}\text{[Hz]}$ #Ch Size [GB] Placement
Ninapro DB6 [ninapro_db6] 10 2000 14 20.3 Forearm
Ninapro DB7 [ninapro_db6] 22 2000 12 30.9 Forearm
EMG2Pose [salter_emg2pose_2024] 192 2000 16 431 Wrist
Ninapro DB6 [ninapro_db6] contains sEMG from 10 subjects performing 7 hand grasp types, each repeated 12 times over 5 separate recording days. Signals were acquired with Trigno sensors (14 channels, $2~\text{KHz}$ sampling rate), placed on the forearm. The total dataset size is $20.3~\text{GB}$.
Ninapro DB7[ninapro_db7] contains data from 20 intact subjects and 2 transradial amputees, performing 40 movements spanning isolated finger/wrist actions and grasp patterns. Data were collected using Trigno sensors (12 channels, $2~\text{kHz}$ sampling rate), placed on the forearm. The total dataset size is $30.9~\text{GB}$.
EMG2Pose[salter_emg2pose_2024] contains data from 193 participants, performing 29 different hand motions. Data were recorded with a custom wristband (16 channels, $2~\text{kHz}$ sampling rate). The total dataset size is $431~\text{GB}$.
@@ 3.1.5 Pre-training Data Processing
EMG data are pre-processed using a Butterworth band-pass filter ($4^{th}$ order, $20-450~\text{Hz}$), followed by a notch filter ($50~\text{Hz}$), as in [chen_2025_physiowave]. Min-Max, channel-wise normalization ([1,1][-1,1] range) is applied. Each recording is then segmented into fixed-length windows of 1000 samples ($500~\text{ms}$ at $2~\text{kHz}$) with 50% overlap, providing sufficient temporal context [smith_emgwindow_2011][delafuente_emgwindow_2021] while controlling computational load.
@@
We did not apply any data augmentation in line with previous works [chen_2025_physiowave, chen_2025_waveformer] which already showed SoA performances. Additionally, given the small model size (3.6 M parameters) and the limited effect of data augmentation shown in scaling-laws regimes [kaplan2020scalinglawsneurallanguage], we rely on the sample-efficiency of the architecture rather than increasing artificially the data variance.
@@ 3.2.1 Pipeline Overview
For all downstream tasks, the Encoder architecture described in Sec. 3.1.2 is retained without modification, while the Decoder used during pretraining is removed.
@@ 3.2.2 Downstream Tasks
Table 3: Summary of downstream EMG datasets.
Dataset #Subj. $f_{s}[\text{Hz}]$ Ch Size [GB] Placement Window size (ms) Pre-Processing Metrics
Ninapro DB5 [ninapro_db5] 10 200 16 0.2 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1, AUROC
EPN-612 [benalcazar_emg-epn-612_2020] 612 200 8 5.5 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1, AUROC
UC Irvine [emg_data_for_gestures_481] 36 200 8 0.32 Forearm 200, 25% overlap $||$ 1000, 25% overlap BP 20–90 Hz $\rightarrow$ NC 50 Hz Accuracy, F1
Generic Neuromotor Interface [generic_neuromotor_interface_2025] 100 2000 16 31.1 Hand/Arm 8000 (no overlap) HP ($40\text{Hz}$) CLER
Ninapro DB8 [ninapro_db8] 12 2000 16 23.6 Forearm 200 MAE
Gaddy Silent Speech [gaddy-klein-2020-digital] 1 1000 8 3.9 Face/Neck $\approx$3000 Series of NC ($60,...,420]\text{Hz} \rightarrow$ HP $2\text{Hz}$ $\rightarrow$ downsampling WER
BP: band-pass filter; NC: notch filter.
@@
  • Ninapro DB5 [ninapro_db5] contains data collected from 10 intact subjects, using two Thalmic Myo Armbands (16 channels, $200~\text{Hz}$). The task consists of classifying 52 hand-wrist movements.
  • EPN-612 [benalcazar_emg-epn-612_2020] contains data collected from 612 subjects, using the Myo Armband (8 channels, $200~\text{Hz}$). The task consists of classifying 5 hand movements.
  • UC Irvine [emg_data_for_gestures_481] contains data from 36 subjects (8 channels, $200~\text{Hz}$). The task consists of classifying 7 movements.
Datasets are pre-processed using a band-pass filter (Butterworth, $4^{th}$ order, $20-90]~\text{Hz}$), followed by a notch filter ($50~\text{Hz}$). Then, windowing is applied. We perform an ablation over two different window sizes: $200~\text{ms}$ with 25% overlap, and $1000~\text{ms}$ with 25% overlap. After filtering and windowing, each dataset is normalized using per-channel z-score normalization.
Compared to the pre-training, datasets with fewer than 16 channels are not zero-padded because tokenization is performed on a per-channel basis; thus, having fewer channels simply results in fewer tokens, eliminating the need for padding.
For Ninapro DB5, we perform a per-repetition split as in [chen_2025_physiowave], each recording contains six repetitions, and for each subject, we assign whole repetitions to a single split so that no subject data overlaps across splits. Specifically, repetitions 1, 3, 4, and 6 are used for training, repetition 2 for validation, and repetition 5 for testing.
The UCI EMG dataset is split by subject index into three disjoint groups: subjects 1–24 are used for training, subjects 25–30 for validation, and subjects 31–36 for testing. Only subjects 11 and 30 performed the ”extended palm” gesture, so they were excluded from all data splits, hence reducing the total number of gestures to 6.
For EPN-612, we preserve the original per-user organization [benalcazar_emg-epn-612_2020] and derive splits per user. When constructing the testing split, held-out examples per user are used, ensuring that samples from the same user do not appear across train, validation, or test splits.
The model head architecture is lightweight and consists of a single fully connected linear layer that maps the pooled feature representation (max $192\times N_{c}$, $N_{c}=$ number of channels ) to the final class logits ($N_{o}$, $N_{o}=$number of outputs). No additional hidden layers, normalization modules, or non-linearities are used in the head. Hence, the model head has a maximum of $192\times N_{c}\times N_{o}$ parameters. For training, we employ the cross-entropy loss computed over the predicted logits and the corresponding ground-truth labels. We evaluate performance using Accuracy, F1-score, and AUROC, with the respective definitions given below:
Accuracy (micro) $=\frac{1}{N}\sum_{i=1}^{N}1(\hat{y_{i}}=y_{i}),$
Accuracy (macro) $=\frac{1}{K}\sum_{k=1}^{K}\frac{\mathrm{TP}_{k}+\mathrm{TN}_{k}}{\mathrm{TP}_{k}+\mathrm{TN}_{k}+\mathrm{FP}_{k}+\mathrm{FN}_{k}},$
F1-score (macro) $=\frac{1}{K}\sum_{k=1}^{K}\frac{2\,\mathrm{TP}_{k}}{2\,\mathrm{TP}_{k}+\mathrm{FP}_{k}+\mathrm{FN}_{k}},$
AUROC (macro) $=\frac{1}{K}\sum_{k=1}^{K}\mathrm{AUROC}_{k}.$
where $\hat{y}_{i}$ is the predicted class, $y_{i}$ the ground-truth label, and $\mathrm{TP}_{k}$, $\mathrm{FP}_{k}$, $\mathrm{FN}_{k}$ denote true positives, false positives, and false negatives for class $k$.
@@
  • Generic Neuromotor Interface [generic_neuromotor_interface_2025] contains data collected from 4900 participants, recorded with a custom wrist-band developed by Meta Reality Labs (16 channels, $2\text{kHz}$). Data from only 100 subjects are publicly available [metagit], randomly selected from the entire cohort.
Specifically, we tackle the discrete-gesture task presented in the original work, consisting of the classification of 9 fine finger movements. We treat this task separately, compared to the three other datasets presented above, owing to substantial methodological differences in data processing and evaluation protocols.
@@
We evaluate the performance with the Classification Error Rate (CLER), defined as the proportion of ground-truth labels for which the matching prediction is incorrect [generic_neuromotor_interface_2025]. The CLER is computed per-gesture and reported as the average across gestures. The CLER formula is given by:
@@
This task consists of predicting continuous hand and wrist movement (regression task). We employ the Ninapro DB8 [ninapro_db8] dataset. It contains EMG data collected from 12 subjects (16 channels, $2~\text{kHz}$), with 3 sessions per subject. Data were collected with an instrumented glove from ground-truth. The task consists of regressing 5 Degrees of Freedom (DoFs), defined as linear combinations of the 18 DoFs of the instrumented glove [krasoulis_2019_effect].
@@
Each EMG channel is independently standardized using z-score normalization. No filtering is applied since the original data is already released filtered ($4\textsuperscript{th}$ order bandpass Butterworth filter, $10-500~\text{Hz}]$. The glove data undergo a mapping from 18 Degrees of Freedom (DoF) to 5 Degrees of Actuation (DoA). Both EMG and glove data are then segmented into fixed-size windows with no overlap in two different settings: $1000\text{ms}$ and $200\text{ms}$.
@@
Following the author’s recommendations, we consider the first two sessions for training, the third as validation. The model head architecture (788 K parameters) consists of a compact series of pointwise and depthwise convolutions, followed by upsampling to a fixed length. We employ the L1 loss function.
We report the mean absolute error (MAE), root mean squared error (RMSE), and the coefficient of determination $R^{2}$, defined as:
$\mathrm{MAE}=\frac{1}{N}\sum_{i=1}^{N}\lvert y_{i}-\hat{y}_{i}\rvert$.
$\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}.$
$R^{2}=1-\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}.$
where $y_{i}$ are the true values, $\hat{y}_{i}$ the predicted values, and $\bar{y}=\frac{1}{N}\sum_{i=1}^{N}y_{i}$. We conduct 5 runs, with 5 random seed values, and report the average.
Downstream 3: Speech Synthesis and Recognition
@@
We employ the open-vocabulary dataset presented by Gaddy et al. [gaddy-klein-2020-digital]. It comprises 8 channels, sampled at $1000~\text{Hz}$, recorded using wet electrodes, placed on the face. Data were collected from one subject in multiple sessions while reading pieces of public domain books. The dataset contains both vocalized (i.e., normal articulation with sound production) and silent (i.e., articulation without sound production) EMG data. In the vocalized setting, EMG data were recoded together with a microphone, serving as ground truth. For the silent setting, audio recorded during vocalized experiments is aligned to the corresponding silent EMG data using the dynamic time warping algorithm [gaddy-klein-2020-digital]. The publicly available dataset contains already-aligned data.
@@
For both tasks, we adopt the same data processing pipeline as in [gaddy2021improved]. A cascade of notch filters is applied to suppress the first six harmonics of power line frequency ($60,240,...,420]~\text{Hz}$). Additionally, a high-pass filter (Butterworth, $2~\text{Hz}$, 3rd order) is applied. Data are downsampled (from $1000~\text{Hz}$ to $516.79~\text{Hz}$, matching the processing pipeline in [gaddygit]).
@@
For this dataset, we tackle two sub-tasks: Speech Synthesis and Speech Recognition. Speech Synthesis aims at producing audio from EMG data. The full speech production pipeline consists of three steps: a Transduction Model, a Vocoder, and an Automatic Speech Recognition (ASR) model. First, a Transduction model is trained to reconstruct audio target features (Mel-Frequency Cepstral Coefficients, MFCC). It consists of 3 residual blocks (as in [gaddy2021improved]) that downsample the input (1600 samples to 200 samples). The 200 samples are the input to the pre-trained Transformer Encoder; we add relative positional embeddings, following [gaddy2021improved]. A lightweight linear head follows, with an input dimension of $192\times N_{c}$, output dimension of 26 MFCC.
@@
The model is trained using as loss function the mean of pair-wise MFCC distances ($L$) plus an auxiliary phoneme loss ($L_{phon}$)  [gaddy2021improved]. In the case of silent EMG data, DTW is used to align the MFCC features predicted from silent EMG to the sequence of target MFCC features obtained from a parallel vocalized recording. Predicted MFCC audio features are fed to a HiFi-GAN [kong_2020_hifigangenerativeadversarialnetworks] vocoder, which converts the audio features into raw audio waveforms. We adopt the publicly available [gaddygit] pre-trained HiFi-GAN vocoder from Gaddy’s work [gaddy-klein-2020-digital], without any fine-tuning.
@@
For model evaluation, we employ the Wav2Vec2 [baevski_2020_wav2vec20frameworkselfsupervised] ASR model, pretrained on LibriSpeech (English Language) within SpeechBrain [speechbrain].
@@
Speech Recognition consists of recognizing speech (words) without any acoustic output, yielding text as output. Producing text directly from EMG signals is preferable to first synthesizing audio and then applying an ASR system: direct text prediction serves as a more efficient surrogate for human intelligibility, avoids the intermediate audio-generation stage, and yields improved recognition accuracy [gaddy_2022_phdthesis]. Moreover, EMG-to-text models do not require an additional vocoder for speech production, substantially reducing model size and computational load. This makes the approach more suitable for deployment on resource-constrained embedded platforms.
@@
The model architecture is the same as before (Residuals, pre-trained Transformer Encoder and linear layer with softmax). The input to the final linear layers has a dimension of $N_{c}\times 192$, the output dimension is 37 (26 lowercase English letters, 10 digits, and space character). We adopt the Connectionist Temporal Classification (CTC) [graves_2006_connectionist] loss function with a 4-gram language model trained on the LibriSpeech dataset [LibriSpeech]. During inference, beam search is applied to explore the search space and select the paths with the highest likelihood.
@@
The dataset is split as in Gaddy et al. [gaddy2021improved]. Training is performed using both silent and vocalized EMG recordings; testing is performed on silent data. Data are split into training, validation, and test sets using the original index file containing pairs of book and sentence indexes. Each utterance is assigned to one of the three splits. During training, a size-aware sampler is used to group examples, together producing compact batches and a more efficient training stage.
For both tasks, we use the Word Error Rate (WER) as the evaluation metric, defined as
$\text{WER}=\frac{S+D+I}{N},$
@@
where $S$, $D$, and $I$ denote the number of substitution, deletion, and insertion errors, respectively, and $N$ is the total number of words in the reference transcription.

### v2

@@ 3.1.4 Pre-training Datasets
Ninapro DB6 [38] contains sEMG from 10 subjects performing 7 hand grasp types, each repeated 12 times over 5 separate recording days. Signals were acquired with Trigno sensors (14 channels, $2~\text{KHz}$ sampling rate), placed on the forearm. The total dataset size is $20.3~\text{GB}$.
Ninapro DB7[30] contains data from 20 intact subjects and 2 transradial amputees, performing 40 movements spanning isolated finger/wrist actions and grasp patterns. Data were collected using Trigno sensors (12 channels, $2~\text{kHz}$ sampling rate), placed on the forearm. The total dataset size is $30.9~\text{GB}$.
EMG2Pose[43] contains data from 193 participants, performing 29 different hand motions. Data were recorded with a custom wristband (16 channels, $2~\text{kHz}$ sampling rate). The total dataset size is $431~\text{GB}$.
@@ 3.1.5 Pre-training Data Processing
EMG data are band-pass filtered ($4^{th}$ order Butterworth, $20-450~\text{Hz}$), notch filtered ($50~\text{Hz}$), normalized to [1,1][-1,1], and segmented into 1000-sample windows ($500~\text{ms}$ at $2~\text{kHz}$) with 50% overlap, providing sufficient temporal context [46] while controlling computational load.
@@
We did not apply any data augmentation due to the limited effect in scaling-laws regimes [27] and in line with previous works [10, 11].
@@ 3.2.1 Pipeline Overview
For all downstream tasks, the pre-trained encoder is retained and the decoder removed.
@@ 3.2.2 Downstream Tasks
Table 1: Summary of Downstream Datasets.
Dataset #S Hz Ch Loc. Metric
Hand Gesture Classification
Ninapro DB5 [14] 10 200 16 FA Acc, F1
EPN-612 [15] 612 200 8 FA Acc, F1
UCI-EMG [16] 36 200 8 FA Acc, F1
GNI [5] 100 2k 16 H/A CLER
Hand Kinematic Regression
Ninapro DB8 [17] 12 2k 16 FA MAE
Speech Production & Recognition
Gaddy [18] 1 1k 8 F/N WER
Loc: Location (FA: Forearm, H/A: Hand/Arm, F/N: Face/Neck).
Pre-processing and windowing details are provided in Section III-B.
@@
  • Ninapro DB5 [39]: data from 10 intact subjects, using two Thalmic Myo Armbands (16 channels, $200~\text{Hz}$). The task consists of classifying 52 hand-wrist movements.
  • EPN-612 [6] data from 612 subjects, using the Myo Armband (8 channels, $200~\text{Hz}$). The task consists of classifying 5 hand movements.
  • UC Irvine [32] data from 36 subjects (8 channels, $200~\text{Hz}$). The task consists of classifying 7 movements.
Datasets are pre-processed using a band-pass filter (Butterworth, $4^{th}$ order, $20-90]~\text{Hz}$), followed by a notch filter ($50~\text{Hz}$). Then, windowing is applied. We perform an ablation over two different window sizes: $200~\text{ms}$ with 25% overlap, and $1000~\text{ms}$ with 25% overlap. Finally, each dataset is normalized using per-channel z-score normalization.
Compared to the pre-training, datasets with fewer than 16 channels are not zero-padded because tokenization is performed on a per-channel basis: having fewer channels simply results in fewer tokens, eliminating the need for padding.
For Ninapro DB5, we perform a per-repetition split as in [10], each recording contains six repetitions, and for each subject, we assign whole repetitions to a single split so that no subject data overlaps across splits (repetitions 1-3-4-6 are used for training, rep. 2 for validation, and rep. 5 for testing).
The UCI EMG dataset is split by subject index into three disjoint groups: subjects 1–24 for training, subjects 25–30 for validation, and subjects 31–36 for testing. Only subjects 11 and 30 performed the ”extended palm” gesture, so they were excluded from all data splits, hence reducing the total number of gestures to 6.
For EPN-612, we preserve the original per-user organization [6] and derive splits per user. When constructing the testing split, held-out examples per user are used, ensuring that samples from the same user do not appear across train, validation, or test splits.
The model head architecture is lightweight and consists of a single fully connected linear layer that maps the pooled feature representation (max $192\times N_{c}$, $N_{c}=$ number of channels ) to the final class logits ($N_{o}$, $N_{o}=$number of outputs). No additional hidden layers, normalization modules, or non-linearities are used in the head. Hence, the model head has a maximum of $192\times N_{c}\times N_{o}$ parameters. For training, we employ the cross-entropy loss computed over the predicted logits and the corresponding ground-truth labels. We evaluate performance using Accuracy, F1-score, and AUROC.
@@
  • Generic Neuromotor Interface [26] contains data collected from 4900 participants, recorded with a custom wrist-band developed by Meta Reality Labs (16 channels, $2\text{kHz}$). Data from only 100 subjects are publicly available [1], randomly selected from the entire cohort.
Specifically, we tackle the discrete-gesture task presented in the original work, i.e., classifying 9 fine finger movements. We treat this task separately, compared to the three other datasets presented above, owing to substantial methodological differences in data processing and evaluation protocols.
@@
We evaluate the performance with the Classification Error Rate (CLER), defined as the proportion of ground-truth labels for which the matching prediction is incorrect [26]. The CLER is computed per-gesture and reported as the average across gestures. The CLER formula is given by:
@@
This task consists of predicting continuous hand and wrist movement (regression). We employ the Ninapro DB8 [31] dataset, containing EMG data from 12 subjects (16 channels, $2~\text{kHz}$), with 3 sessions per subject. Data were collected with an instrumented glove from ground-truth. The task consists of regressing 5 Degrees of Freedom (DoFs), defined as linear combinations of the 18 DoFs of the instrumented glove.
@@
Each EMG channel is independently standardized using z-score normalization. No filtering is applied since the original data is already released filtered ($4\textsuperscript{th}$ order bandpass Butterworth filter, $10-500~\text{Hz}]$. The glove data undergo a mapping from 18 DoFs to 5 Degrees of Actuation. Both EMG and glove data are then segmented into fixed-size windows with no overlap in two different settings: $1000\text{ms}$ and $200\text{ms}$.
@@
Following the author’s recommendations, we consider the first two sessions for training, the third as validation. The model head architecture (788 K parameters) consists of a compact series of pointwise and depthwise convolutions, followed by upsampling to a fixed length. We employ the L1 loss function. We report the mean absolute error (MAE), root mean squared error (RMSE), and the coefficient of determination $R^{2}$, averaged over 5 runs with 5 random seed values. Downstream 3: Speech Synthesis and Recognition
@@
We employ the open-vocabulary dataset presented by Gaddy et al. [16]. It comprises 8 channels, sampled at $1000~\text{Hz}$, recorded using wet electrodes, placed on the face. Data were collected from one subject in multiple sessions while reading pieces of public domain books. The dataset contains both vocalized (i.e., normal articulation with sound production) and silent (i.e., articulation without sound production) EMG data. In the vocalized setting, EMG data were recoded together with a microphone, serving as ground truth. For the silent setting, audio recorded during vocalized experiments is aligned to the corresponding silent EMG data using the dynamic time warping algorithm [16].
@@
For both tasks, we adopt the data processing pipeline of [17]. A cascade of notch filters is applied to suppress the first six harmonics of power line frequency ($60,240,...,420]~\text{Hz}$), followed by a high-pass filter (Butterworth, $2~\text{Hz}$, 3rd order) and downsampling from to $516.79~\text{Hz}$ [51].
@@
For this dataset, we tackle two sub-tasks: Speech Synthesis and Speech Recognition. Speech Synthesis aims at producing audio from EMG data. The full speech production pipeline consists of three steps: a Transduction Model, a Vocoder, and an Automatic Speech Recognition (ASR) model. First, a Transduction model is trained to reconstruct audio target features (Mel-Frequency Cepstral Coefficients, MFCC). It consists of 3 residual blocks (as in [17]) that downsample the input (1600 samples to 200 samples). The 200 samples are the input to the pre-trained Transformer Encoder; we add relative positional embeddings, following [17]. A lightweight linear head follows, with an input dimension of $192\times N_{c}$, output dimension of 26 MFCC.
@@
The model is trained using as loss function the mean of pair-wise MFCC distances ($L$) plus an auxiliary phoneme loss ($L_{phon}$)  [17]. For silent EMG data, DTW is used to align the MFCC features predicted from silent EMG to the sequence of target MFCC features obtained from a parallel vocalized recording. Predicted MFCC audio features are fed to a HiFi-GAN vocoder, which converts the audio features into raw audio waveforms. We adopt the publicly available pre-trained HiFi-GAN vocoder [51] without any fine-tuning.
@@
For model evaluation, we employ the Wav2Vec2 [4] ASR model, pretrained on LibriSpeech (English Language) within SpeechBrain [41].
@@
Speech Recognition consists of recognizing speech (words) without any acoustic output, yielding text as output. Direct EMG-to-text prediction avoids intermediate audio synthesis and reduces model complexity, yielding improved recognition accuracy [18]. This makes the approach more suitable for deployment on resource-constrained embedded platforms.
@@
The model architecture is the same as before. The input to the final linear layers has a dimension of $N_{c}\times 192$, the output dimension is 37 (26 lowercase English letters, 10 digits, and space character). We adopt the Connectionist Temporal Classification (CTC) [20] loss function with a 4-gram language model trained on the LibriSpeech dataset [34]. During inference, beam search is applied to explore the search space and select the paths with the highest likelihood.
@@
The dataset is split as in [17]. Training is performed using both silent and vocalized EMG recordings; testing is performed on silent data. Data are split into training, validation, and test sets using the original index file containing pairs of book and sentence indexes. Each utterance is assigned to one of the three splits. During training, a size-aware sampler is used to group examples, together producing compact batches and a more efficient training stage.
@@
For both tasks, we use the Word Error Rate (WER) as the evaluation metric, defined as $\text{WER}=\frac{S+D+I}{N}$, where $S$, $D$, and $I$ denote the number of substitution, deletion, and insertion errors, respectively, and $N$ is the total number of words in the reference transcription.